OpenStudy (anonymous):

How can I factor this? 2sin2(squared)β-1=0 ?

OpenStudy (welshfella):

use difference of 2 squares

OpenStudy (welshfella):

a^2 - b^2 = (a + b)(a - b)

OpenStudy (welshfella):

in you case put a = sqrt2 sinβ and b = 1

OpenStudy (er.mohd.amir):

cos 2x =1-2 sin ^2 x

OpenStudy (anonymous):

Like this? \[2 \sin \beta - \sin \beta - 1 \]

OpenStudy (welshfella):

if you just want to solve the equation then Mohd's method is more straightforward

OpenStudy (welshfella):

no factorisung would give (sqrt2 sin β + 1)(sqrt2 sinβ - 1) = 0

OpenStudy (welshfella):

this will give sinβ = 1/sqrt2 or -1/sqrt2

OpenStudy (welshfella):

so there are 3 solution in the range 0 to 360 degrees 1 in the first quadrant , 1 in the 3rd and 1 in the 4th.